OpenStudy (anonymous):

How long will it take a $500 investment to be worth $900 if it is continuously compounded at 12% per year? - I keep getting 7.05 years, but that's not the right answer!

OpenStudy (anonymous):

simple

OpenStudy (anonymous):

if it's compounded annually A=P(1+r)^t 900=500(1.12)^t t=5.187

OpenStudy (anonymous):